### v1

@@ 3 Preliminaries
System model. $\mathsf{ammBoost}$ involves a base smart contract representing the AMM on the mainchain, and a sidechain that processes most of the AMM workload. Anyone can join/leave the AMM at anytime, and these parties are known using their public keys. Participants are three types: clients $\mathcal{C}$ who are only interested in using the AMM trading services, liquidity providers $\mathcal{LP}$ who provide liquidity for the pools operated by the AMM, and miners $\mathcal{M}$ who maintain the AMM sidechain. We do not place any restrictions on the mainchain beyond being a secure smart contract-enabled blockchain. $\mathsf{ammBoost}$ operates in rounds and epochs (as defined earlier). The sidechain is managed by a committee elected from the sidechain miners, where a new committee is elected for each epoch. This committeee runs a PBFT-based consensus to mine new blocks. These blocks are as in the chainBoost framework: temporary meta-blocks that record transactions, and permanent summary-blocks that summarize meta-blocks mined in an epoch. The committee also issues sync-transactions to sync the base AMM smart contract deployed on the mainchain.
Accordingly, the $\mathsf{ammBoost}$ framework provides the following functionalities:
@@
Takes as input transaction type $\mathsf{txtype}$ and any additional information/inputs $\mathsf{aux}$, and outputs a transaction $\mathsf{tx}$ of one of the following types:
  • $\mathsf{tx}_{\mathsf{Deposit}}$: Allows a user (client or LP) to deposit funds on the mainchain to support her activities on the sidechain.
@@
  • $\mathsf{tx}_{\mathsf{mint}}$: Allows an LP to provide liquidity to a particular pool.
@@
  • $\mathsf{tx}_{\mathsf{burn}}$: Allows an LP to withdraw her liquidity position.
  • $\mathsf{tx}_{\mathsf{Sync}}$: Allows a sidechain committee to sync the AMM base contract on the mainchain.
@@
Takes as input the current sidechain ledger state $\mathcal{L}_{\mathsf{sc}}$, a new block $\mathsf{B}$ with type $\mathsf{btype}=\mathsf{meta}$ or $\mathsf{btype}=\mathsf{summary}$. It outputs 1 if $\mathsf{B}$ is valid based on the syntax/semantics of the particular block type, and 0 otherwise.
@@
Takes as input the current sidechain state $\mathcal{L}_{\mathsf{sc}}$, and a set of pending transactions $\mathsf{aux}=\{\mathsf{tx}_{i}\}$ (if $\mathsf{btype}=\mathsf{meta}$) or $\bot$ (if $\mathsf{btype}=\mathsf{summary}$ since the inputs are the last epoch meta-blocks from $\mathcal{L}_{\mathsf{sc}}$). It reflects the changes induced by $\mathsf{aux}$ and outputs a new ledger state $\mathcal{L}_{\mathsf{sc}}^{\prime}$.
@@
Takes as input the current sidechain ledger state $\mathcal{L}_{\mathsf{sc}}$, and produces an updated state $\mathcal{L}_{\mathsf{sc}}^{\prime}$ in which all stale meta-blocks are dropped.
Note that $\mathsf{UpdateState}$ is the process of mining a new block on the sidechain. Agreement on this block is governed by the sidehchain consensus protocol.
@@
Adversary model. Our model is very similar to the one in [54]. In particular, we assume the mainchain to be secure as defined above. For the sidechain, we have honest miners who follow the protocol, and malicious miners controlled by the adversary who may behave arbitrarily. The adversary can deploy new miners or corrupt existing ones, without going above the threshold of faulty nodes required by the sidechain consensus protocol. The adversary can see all messages and transactions sent in the system (since we deal with public open-access blockchain systems) and can reorder these messages and delay them. We assume bounded-delay message delivery, so any sent message (or transaction) will be delivered within $\Delta$ time as in [47, 56, 51]. We assume slowly-adaptive adversaries [33] that can corrupt miners only at the beginning of each epoch. Lastly, we deal with $\mathsf{PPT}$ adversaries.
@@ 4 System Design
$\mathsf{ammBoost}$ changes the AMM deployment structure, as shown in Figure 2. The smart contract on the mainchain will become minimal; it mainly tracks the token balances of the liquidity pools and the users, while most of the transaction processing is moved to the sidechain. Summaries of the sidechain processed traffic are used to sync the AMM smart contract on the mainchain. Also, the AMM now has its own miners to maintain the sidechain. In this section, we present the design of $\mathsf{ammBoost}$ including system setup, architecture and operation, handling interruptions, and its security.
@@ 4.1 System Setup
The setup phase, as depicted in Figure 3, mainly specifies the traffic split between the chains and the summary rules for the sidechain traffic. Also, this phase involves deploying the AMM base smart contract on the mainchain and creating the sidechain.
Traffic classification and summary rules. AMM transactions and operations are divided into two groups: pool management and trading-related operations. Creating and managing token pools, as well as dispensing tokens to clients and LPs, are done on the mainchain (using the $\mathsf{TokenBank}$ contract as we explain shortly). Flashes are also processed on the mainchain since they require instant token dispensing rather than at the end of the epoch (which is the case for any operation done on the sidechain). The rest of the transactions, including swaps, mints, burns, and collects, are handled by the sidechain.
In $\mathsf{ammBoost}$, the sidechain does not store any real tokens, it just tracks their balances and update them based on the transactions this sidechain processes. That is, during each epoch, two structures are produced:
  • A payout list containing users’ public keys and the amount/type of tokens they should receive. These result from swaps, withdrawal of liquidity positions, and fee collection.
  • A payin list containing also the users’ public keys and the amount/type of tokens that should be deducted from their mainchain deposits for their sidechain activities. These result from swaps and mints.
$\mathsf{SystemSetup}(1^{\lambda},\mathcal{L}_{\mathsf{mc}})$: Takes as input the security parameter $\lambda$ and the current state of the mainchain $\mathcal{L}_{\mathsf{mc}}$, and does the following:
@@ 1. Generate the sidechain configuration parameters:
2. Deploy the AMM base smart contract $\mathsf{TokenBank}$ on the mainchain.
@@
The actual token dispensing and deduction happens at the end of an epoch when the sidechain summaries are received. The new state of the pool token balances on the mainchain will be computed based on these lists. Also, tokens will be sent to users (i.e., payouts) or added to the pool from the users’ deposits (i.e., payins). In Section 4.2, we show the summary rules for each transaction type and how they contribute to the payout and payin lists.
@@
$\mathsf{Deposit}(\mathsf{type},\mathsf{amnt})$: allows a user to deposit an amount $\mathsf{amnt}$ of token with type $\mathsf{type}$ to back up her activities for the next epoch.
$\mathsf{Sync}(\mathsf{aux})$: Sync the on-chain AMM state based on the sidechain epoch summaries. The input $\mathsf{aux}$ contains the updated balances of the pools, updated liquidity positions, and the payin/payout lists.
@@
Base smart contract $\mathsf{TokenBank}$. The mainchain part of the AMM is a base smart contract called $\mathsf{TokenBank}$. At an abstract level, and as shown in Figure 4, this contract supports creating and managing token pools (the latter involves tracking the token balances in these pools and their liquidity positions). It also provides the minimal interface needed to support users’ activities on the sidechain, which is mainly creating deposits on the mainchain of the tokens they want to trade or provide as liquidity. This is needed since the sidechain does not receive or send actual tokens, but rather it tracks their state evolution. Hence, a user deposits the total amount of tokens, in the $\mathsf{TokenBank}$ account, that is needed to support their activities during an epoch before this epoch starts. This also means that $\mathsf{TokenBank}$ handles payouts and payins that the sidechain produces at the end of any epoch.
As shown in Figure 3, system designers create and deploy the $\mathsf{TokenBank}$ contract on the mainchain. Once the mainchain block containing this contract is confirmed, the genesis block of the sidechain $\mathcal{L}_{\mathsf{sc}}^{0}$ can be created such that it references this mainchain block.
@@
$\mathsf{ammBoost}$ differs from chainBoost in the aspect that the sidechain has its own miner population. In other words, mainchain miners are not responsible of managing the sidechain, and even may not know that a sidechain exists in the first place. Executing the $\mathsf{TokenBank}$ contract is like executing any other contract deployed at the application layer of the mainchain. As such, sidechain miners must posses some mining power to establish Sybil resistant identities to be used in the committee election process. Any secure PBFT protocol in which election is based on the mining power can be used here, e.g., the proof-of-stake based protocol in [47].
4.2 System Operation
As mentioned before, $\mathsf{ammBoost}$ operates in epochs and rounds. The sidechain committee begins the epoch by retrieving the latest state, i.e., pool token balances, liquidity positions, and the deposit balances of users (both clients and LPs) from the mainchain. It then processes all valid sidechain transactions, including swap, mint, burn, and collect. These are packaged into meta-blocks such that a meta-block is mined in each round. In the last round of the epoch, this committee produces a summary-block capturing the payouts and payins for participating users, any changes on liquidity positions’ ownership and amounts, and the updated liquidity pool balance. After that, it invokes the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles submitting a sync-transaction to update the AMM state on the mainchain, which is basically the state of the $\mathsf{TokenBank}$ contract.
In what follows, we describe how the various transactions are processed and summarized (Figure 5 captures how the sidechain workload is summarized in $\mathsf{ammBoost}$).
Swaps. A swap transaction is a trade between the two tokens managed by a liquidity pool. A client provides an input of tokens and receives an output of the other token type with an amount based on the price derived from the pool token balances and any user-defined trade conditions.
Input: meta-blocks $\mathsf{B}^{1}_{\mathsf{meta}},\dots,\mathsf{B}^{n}_{\mathsf{meta}}$ from an epoch.
Initialize empty summary structures $\mathsf{sum}_{\mathsf{Payins}}$, $\mathsf{sum}_{\mathsf{Payouts}}$, and $\mathsf{sum}_{\mathsf{Positions}}$
@@ 4.2 System Operation
$\mathsf{sum}_{\mathsf{Payins}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{in}.\mathsf{type}]\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{\mathsf{in}}$
$\mathsf{sum}_{\mathsf{Payouts}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{out}.\mathsf{type}]\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{\mathsf{out}}$
@@
// Liquidity amounts are computed as explained under mints and burns.
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{priceRange}=(\mathsf{tx}.\mathsf{lowerTick},\mathsf{tx}.\mathsf{upperTick})$
$\mathsf{sum}_{\mathsf{Payins}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Payins}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{sum}_{\mathsf{Payouts}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Payouts}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{sum}_{\mathsf{Payouts}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{sum}_{\mathsf{Payouts}}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
Output $\mathsf{sum}_{\mathsf{Payins}}$, $\mathsf{sum}_{\mathsf{Payouts}}$, and $\mathsf{sum}_{\mathsf{Positions}}$
@@
Figure 5: Summary rules ($\mathsf{userId}$ is the user ID and $\mathsf{posId}$ is the liquidity position ID). Updated liquidity pool balances are computed by TokenBank (as part of processing $\mathsf{Sync}$) based on the updated liquidity position list.
@@
In order to execute a swap transaction, a user must have already deposited their input tokens in the $\mathsf{TokenBank}$ contract. Then, they can broadcast their swap transaction to the sidechain. An exact input swap transaction contains the following arguments: type and amount of input tokens to be traded, the minimum amount of output tokens the trade will accept (as a protection against slippage), a price limit that the trade should not exceed, and a deadline which is a round number after which the trade becomes invalid if not executed by that time. [^3] The recipient of the traded tokens is by default the issuer of the swap transaction. This can be extended to allow for stating an explicit recipient that could be different from the issuer. In the case of an exact output swap, the goal is no longer to trade the exact amount of input tokens for the maximum amount of output tokens, but rather to minimize the amount of input tokens required to trade for the desired exact output. As such, the arguments of the function naturally change to reflect that, with the minimum output slippage protection changing to a maximum input slippage protection.
@@
Processing. This is done using the original AMM logic for price balancing and output calculation. That is, $\mathsf{ammBoost}$ does not change the logic based on which an AMM operates, but it rather migrates that process to the sidechain (and this applies to the rest of the transactions as well). Thus, for an exact input swap, the sidechain committee computes the maximum amount of output tokens the user will receive for all of the input tokens provided. While for an exact output swap, the committee computes the minimum amount of input tokens needed to purchase the defined output. In both cases, these computations are based off the updated pool balance on the sidechain. In other words, as transactions are processed, the committee updates the pool state that was retrieved at the beginning of the epoch.
@@
Furthermore, the fees for LPs whose liquidity was used in filling a swap will be computed. To elaborate, when a user submits a swap transaction, they pay a small additional fee, like $0.3\%$ of their transactions input or output value. It is paid in the token pair of the pool based off its net liquidity such that the token with the largest amount of net liquidity is used. For example, if a user provides 100 $A$ input tokens in an exact input swap, and token $A$ is the dominant token, then 0.3 $A$ tokens are used for the LP fee, and 99.7 $A$ tokens is used for the swap transaction. These fees are split up proportionally amongst the positions (based on the amount of liquidity they provide) that occupy the price range for which the swap was executed. $\mathsf{ammBoost}$ maintains a per-position fee balance, which is updated on every swap transaction, again using the same logic used by the underlying AMM to compute these fees.
@@
Lastly, recall that a user’s deposit state is retrieved from $\mathsf{TokenBank}$ at the beginning of each epoch, and swaps (and mints as we will see shortly) will be accepted as long as there is enough deposit to support them. Moreover, a user will not get her actual traded tokens until the end of the epoch. However, this user can use these tokens for trading on the sidechain since the sidechain tracks all balances. So basically, when a swap is executed, the input token amount are deducted from the deposit while the output token amount is added to another deposit balance (maintained on the sidechain) allowing the user to use them immediately for swaps and mints. As a result, the sidechain committee maintains two deposit variables for each user: one for the mainchain deposit and initialized with the value found in $\mathsf{TokenBank}$, and another for accrued sidechain deposit that is initialized to 0 at the beginning of each epoch. Whenever the user issues a swap that uses the sidechain deposit, the amount of input tokens is deducted. This also means that for each swap, the sidechain committee first checks if the mainchain deposit can cover it, and if not checks if the sidechain deposit can cover it and adjust the used deposit balance accordingly.
@@
Summary rules. In the summary-block, the sidechain committee summarizes all swap transactions as follows: for every client of a pool, all of the swaps are combined into a single tuple. The tuple contains the public key of the client, the total payin that should be deducted from her mainchain deposit (i.e., initial mainchain deposit balance minus the updated balance computed on the sidechain), and the total payout that this user should receive, which is simply the updated sidechain deposit.
Mints. Mint transactions consist of either the creation of a new liquidity position, or supplementing additional liquidity into an already existing one.
In order to execute a mint transaction, an LP must begin by first transferring tokens to the $\mathsf{TokenBank}$ contract before the epoch begins so these will be used to provide liquidity as desired. During the epoch, the LP can broadcast a mint transaction that contains the following inputs: the lower and upper ticks, representing the price range for which the liquidity is to reside in, and the type/amount of the token to be used as liquidity.
@@
Processing. This is also processed using the same logic used by the AMM. We resort to a simple approach to track ownership of positions; the sidechain committee generates a unique identifier (e.g., the hash of the mint transaction along with the LP’s public key) for a new position, and the owner is the public key of the issuer LP. An existing position will receive an increase in its balance (or any other modifications on its price range) after verifying that the transaction issuer is indeed the rightful owner of the position. The latter is simply done by comparing the public keys of the owner and the transaction issuer. Mint transactions are initially invoked with a desired amount of token $A$ and token $B$ as input. The underlying AMM algorithms compute the maximum amount of liquidity (based off the input tokens) that the pool can take in at the current moment and return that value as an amount of token $A$ and token $B$ (this amount is usually less or equal to the LP’s desired amount). These values represent the share of the pools liquidity now owned by the newly minted or modified position. The mint will be accepted if the issuer LP’s deposit (either the mainchain or sidechain one) can cover the provided liquidity amount (from both tokens managed by the pool), and then this amount will be deducted from the corresponding deposit.

### v3

@@ 3 Preliminaries
System model. $\mathsf{ammBoost}$ involves a base smart contract representing the AMM on the mainchain, and a sidechain that processes most of the AMM workload. Anyone can join/leave the AMM at anytime, and these parties are known using their public keys. Participants are three types: clients $\mathcal{C}$ who are only interested in using the AMM trading services, liquidity providers $\mathcal{LP}$ who provide liquidity for the pools operated by the AMM, and miners $\mathcal{M}$ who maintain the AMM sidechain. We do not place any restrictions on the mainchain beyond being a secure smart contract-enabled blockchain. $\mathsf{ammBoost}$ operates in rounds and epochs (as defined earlier). The sidechain is managed by a committee elected from the sidechain miners, where a new committee is elected for each epoch. This committeee runs a PBFT-based consensus to mine new blocks: temporary meta-blocks that record transactions, and permanent summary-blocks that summarize meta-blocks mined in an epoch. The committee also issues sync-transactions to sync the base AMM smart contract deployed on the mainchain. Accordingly, the $\mathsf{ammBoost}$ framework provides the following functionalities:
@@
Takes the transaction type $\mathsf{txtype}$ and any additional information $\mathsf{aux}$ as inputs, and outputs a transaction $\mathsf{tx}$ of one of the following types:
  • $\mathsf{tx}_{\mathsf{Deposit}}$: Allows a user to deposit funds on the mainchain to support their activities on the sidechain.
@@
  • $\mathsf{tx}_{\mathsf{mint}}$: Allows an LP to provide liquidity to a pool.
@@
  • $\mathsf{tx}_{\mathsf{burn}}$: Allows an LP to withdraw her liquidity.
  • $\mathsf{tx}_{\mathsf{Sync}}$: Allows a sidechain committee to sync the AMM base contract that resides on the mainchain.
@@
Takes as input the current sidechain ledger state $\mathcal{L}_{\mathsf{sc}}$, a new block $\mathsf{B}$ with type $\mathsf{btype}=\mathsf{meta}$ or $\mathsf{btype}=\mathsf{summary}$. It outputs 1 if $\mathsf{B}$ is valid based on the syntax/semantics of the block type, and 0 otherwise.
@@
Takes as input the current sidechain state $\mathcal{L}_{\mathsf{sc}}$, and a set of pending transactions $\mathsf{aux}=\{\mathsf{tx}_{i}\}$ (if $\mathsf{btype}=\mathsf{meta}$) or $\bot$ (if $\mathsf{btype}=\mathsf{summary}$ since the inputs are the last epoch meta-blocks from $\mathcal{L}_{\mathsf{sc}}$). It reflects the changes induced by $\mathsf{aux}$ and outputs a new state $\mathcal{L}_{\mathsf{sc}}^{\prime}$.
@@
Takes as input the current sidechain state $\mathcal{L}_{\mathsf{sc}}$, and produces an updated state $\mathcal{L}_{\mathsf{sc}}^{\prime}$ in which all stale meta-blocks are dropped.
Note that $\mathsf{UpdateState}$ is the process of mining a new block on the sidechain based on its consensus protocol.
@@
Adversary model. We assume the mainchain to be secure as defined above. For the sidechain, we have honest miners who follow the protocol, and malicious miners controlled by the adversary who may behave arbitrarily. The adversary can deploy new miners or corrupt existing ones, without going above the threshold of faulty nodes required by the sidechain consensus protocol. The adversary can see all messages and transactions sent in the system (since we deal with public permissionless blockchain systems) and can reorder these messages and delay them. We assume bounded-delay message delivery, so any sent message (or transaction) will be delivered within $\Delta$ time as in [48, 59, 53]. We assume slowly-adaptive adversaries [34] that can corrupt miners only at the beginning of each epoch. Lastly, we deal with $\mathsf{PPT}$ adversaries.
@@ 4 System Design
$\mathsf{ammBoost}$ changes the AMM deployment structure, as shown in Figure 2. The smart contract on the mainchain is minimal; it mainly tracks the token balances of the liquidity pools and the users, while most of the transaction processing is moved to the sidechain. Summaries of the sidechain processed traffic are used to sync the AMM smart contract on the mainchain. In this section, we present the design of $\mathsf{ammBoost}$ including system setup, architecture and operation, handling interruptions, and its security.
@@ 4.1 System Setup
The setup phase, as depicted in Figure 3, mainly specifies the traffic split between the chains and the summary rules for the sidechain traffic, as well as the sidechain parameters such as the epoch length and its consensus configuration (e.g., committee size). [^7] The epoch duration impacts syncing frequency. Short epochs mean more sync-transactions, which incurs more gas cost and may impact throughput—as they are processed by the mainchain, however users would receive their tokens faster compared to long epochs. We empirically study the impact of this parameter in Section 6. Also, this phase involves deploying the AMM base smart contract on the mainchain and creating the sidechain.
Traffic classification and summary rules. AMM transactions and operations are divided into two groups: pool management and trading-related. Creating and managing token pools, as well as dispensing tokens to clients and LPs, are done on the mainchain since these deal with actual tokens. Flashes are also handled by the mainchain since they require instant token dispensing rather than at the end of the epoch (which is the case for any operation processed by the sidechain). The rest of the transactions, including swaps, mints, burns, and collects, are handled by the sidechain.
In $\mathsf{ammBoost}$, the sidechain does not hold custody of tokens, it just tracks their balances based on the processed transactions. Thus, during each epoch, the sidechain produces two structures:
  • A payout list containing users’ public keys and the amount/type of tokens they should receive. This list is simply the updated deposit balance produced at the end of an epoch.
  • A liquidity position list containing the position IDs, the public keys of their owners, balances, and any additional information needed by the liquidity management techniques, e.g., price ranges as in concentrated liquidity.
$\mathsf{SystemSetup}(1^{\lambda},\mathcal{L}_{\mathsf{mc}})$: Takes as input the security parameter $\lambda$ and the current mainchain state $\mathcal{L}_{\mathsf{mc}}$, and does the following:
@@ 1. Generate the sidechain configuration parameters:
2. Deploy the base contract $\mathsf{TokenBank}$ on the mainchain.
@@
The actual token dispensing and deduction happen at the end of an epoch when the sidechain summaries are received. The new state of the pool token balances on the mainchain will be computed based on these lists. Also, the payout encompasses refunding any leftover in the deposits to their owners as will be shown shortly. In Section 4.2, we show the summary rules for each transaction type and how they contribute to the payout and payin lists.
@@
$\mathsf{Deposit}(\mathsf{type},\mathsf{amnt})$: allows a user to deposit an amount $\mathsf{amnt}$ of token with type $\mathsf{type}$ to be used for the next epoch.
$\mathsf{Sync}(\mathsf{aux})$: Sync the mainchain AMM state with the sidechain epoch summaries. The input $\mathsf{aux}$ contains the updated pool balances and liquidity positions, and the payin/payout lists.
@@
Base smart contract $\mathsf{TokenBank}$. The mainchain part of the AMM is a base smart contract called $\mathsf{TokenBank}$. At an abstract level, as shown in Figure 4, this contract supports creating and managing token pools (i.e., tracking their balances and liquidity positions). It also provides the minimal interface needed to support users’ activities on the sidechain, which is mainly creating deposits containing the tokens they want to trade or provide as liquidity. This is needed since the sidechain does not receive or send actual tokens, it only tracks balance evolution. Hence, a user deposits the total amount of tokens they would need during an epoch before this epoch starts, and $\mathsf{TokenBank}$ handles the payouts and payins produced by the sidechain when the epoch ends.
As shown in Figure 3, system designers deploy $\mathsf{TokenBank}$ on the mainchain. Once the mainchain block containing this contract is confirmed, the genesis block of the sidechain $\mathcal{L}_{\mathsf{sc}}^{0}$ can be created such that it references this mainchain block.
@@
$\mathsf{ammBoost}$ differs from chainBoost in the aspect that the sidechain has its own miner population. In other words, mainchain miners are not responsible for managing the sidechain, and even may not know that a sidechain exists in the first place. Executing the $\mathsf{TokenBank}$ contract is like executing any other contract deployed at the application layer of the mainchain. As such, sidechain miners must possess some mining power to establish Sybil-resistant identities to be used in the committee election process. Any secure PBFT protocol in which election is based on the mining power can be used here, e.g., the proof-of-stake based protocol in [48].
4.2 System Operation—Transaction Processing
As mentioned before, $\mathsf{ammBoost}$ operates in epochs and rounds. The sidechain committee begins the epoch by retrieving the latest state, i.e., pool token balances, liquidity positions, and user deposits from the mainchain. It then processes all valid sidechain transactions, including swap, mint, burn, and collect (so users send these transactions to the sidechain). These are packaged into meta-blocks such that a meta-block is mined in each round. In the last round of the epoch, this committee produces a summary-block capturing the payouts for participating users, and any changes on liquidity positions, where the updated liquidity pool balances will be computed based on these lists. After that, it invokes the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles submitting a sync-transaction to update the AMM state on the mainchain, which is the state of $\mathsf{TokenBank}$.
In this section, we describe how the various transactions are processed and summarized (Figure 5 captures how the sidechain workload is summarized in $\mathsf{ammBoost}$). Before that, we want to point out that although a user will obtain her newly claimed tokens at the end of an epoch, once $\mathsf{TokenBank}$ is synced, they can use these tokens immediately during an epoch for trading. This is because new tokens will be added to the user deposit balance, and those that were used are deducted from deposit. Thus, the latest deposit state reflects the payout a user obtains (which includes refunding any deposit leftover).
Swaps. A swap transaction is a trade between the two tokens managed by a liquidity pool. A client provides an input of tokens and receives an output based on the price derived from the pool token balances and any user-defined trade conditions.
Input: meta-blocks $\mathsf{B}^{1}_{\mathsf{meta}},\dots,\mathsf{B}^{n}_{\mathsf{meta}}$ from an epoch and $\mathsf{Deposits}$ (the latter is the one retrieved from the mainchain at the beginning of the epoch).
Initialize: summary structures $\mathsf{sum}_{\mathsf{Payouts}}$ and $\mathsf{sum}_{\mathsf{Positions}}$.
@@ 4.2 System Operation—Transaction Processing
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{in}.\mathsf{type}]\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{\mathsf{in}}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}[\mathsf{out}.\mathsf{type}]\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{\mathsf{out}}$
@@
// Liquidity amounts are computed as explained under
//mints and burns.
@@
$\mathsf{sum}_{\mathsf{Positions}}[\mathsf{tx}.\mathsf{posId}].\mathsf{priceRange}=$
$(\mathsf{tx}.\mathsf{lowerTick},\mathsf{tx}.\mathsf{upperTick})$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{-}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
@@
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{A}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{A}$
$\mathsf{Deposits}[\mathsf{tx}.\mathsf{userId}].\mathsf{amount}_{B}\mathrel{+}=\mathsf{tx}.\mathsf{amount}_{B}$
Output $\mathsf{sum}_{\mathsf{Payouts}}=\mathsf{Deposits}$, and $\mathsf{sum}_{\mathsf{Positions}}$
@@
Figure 5: Summary rules ($\mathsf{userId}$ is the user ID and $\mathsf{posId}$ is the liquidity position ID). Updated liquidity pool balances are computed by TokenBank (as part of processing $\mathsf{Sync}$) based on the updated liquidity position and payout lists.
@@
In order to execute a swap transaction, a user’s deposit must cover the input token amount. An exact input swap transaction contains: the type and amount of input tokens to be traded, the minimum amount of output tokens the trade will accept (as a protection against slippage), a price limit that the trade should not exceed, and a deadline which is a round number after which the trade becomes invalid if not executed by that time. [^8] The recipient of the traded tokens is by default the issuer of the swap. This can be extended to support stating an explicit recipient that could be different from the issuer. For an exact output swap, the goal is no longer to trade the exact amount of input tokens for the maximum amount of output tokens, but rather to minimize the amount of input tokens required to trade for the desired exact output. As such, the arguments of the function naturally change to reflect that, with the minimum output slippage protection changing to a maximum input slippage protection.
@@
Processing. This is done using the original AMM logic for price balancing and output calculation. That is, $\mathsf{ammBoost}$ does not change the logic based on which an AMM operates, it just migrates that to the sidechain (this applies to the rest of the transactions as well). For an exact input swap, the sidechain committee computes the maximum amount of output tokens the user will receive for all of the input tokens provided. While for an exact output swap, the committee computes the minimum amount of input tokens needed to purchase the defined output. In both cases, these computations are based off the updated pool balance on the sidechain. In other words, as transactions are processed, the committee updates the pool state that was retrieved at the beginning of the epoch.
@@
Furthermore, the fees for LPs whose liquidity was used in filling a swap will be computed. To elaborate, when a user submits a swap transaction, they pay a small additional fee, like $0.3\%$ of their transaction’s input or output value. It is paid in the token pair of the pool based on its net liquidity such that the token with the largest amount of net liquidity is used. For example, if a user provides 100 $A$ input tokens in an exact input swap, and token $A$ is the dominant token, then 0.3 $A$ tokens are used for the LP fee, and 99.7 $A$ tokens are used for the swap transaction. These fees are split up proportionally amongst the positions (based on the amount of liquidity they provide) that occupy the price range for which the swap was executed. $\mathsf{ammBoost}$ maintains a per-position fee balance, which is updated on every swap transaction, again using the same logic used by the underlying AMM to compute these fees.
@@
Lastly, recall that a user will not get her actual traded tokens until the end of the epoch. However, she can use these tokens for trading on the sidechain since the sidechain tracks all balances. So basically, the deposit is a tuple of two values; one for each token type. When a swap is executed, the input token amount is deducted from the user’s deposit while the output token amount is added to this deposit, thus allowing the user to use it immediately.
@@
Summary rules. In the summary-block, the committee summarizes all swap transactions as follows: for every client, all her swaps are combined into a single tuple containing: the client public key, the total payout this user should receive. The latter encapsulates both a deduction from her deposit and a refund of any leftover in that deposit. For example, say a user started with a deposit of $(10A,15B)$ and issued one swap during an epoch that traded 5 token $A$ for 10 token $B$. The updated deposit (which will be the payout summary for that user) would be $(5A,25B)$, which represents a payout of 10 $B$ tokens, a deduction of 5 $A$ tokens, and a refund of deposit leftover of $(5A,15B)$. The same logic applies to the rest of the transactions.
Mints. Mint transactions allow the creation of new liquidity positions or modifying existing ones. An LP broadcasts a mint transaction to the sidechain that contains: the lower and upper ticks, representing the price range for which the liquidity is to reside, and the type/amount of the token to be used as liquidity. The mint will be accepted if the issuer LP’s deposit (either the mainchain or sidechain one) can cover the provided liquidity amount.
@@
Processing. This is also processed using the same logic used by the AMM. We resort to a simple approach to track ownership of positions; the sidechain committee generates a unique identifier (e.g., the hash of the mint transaction and the LP’s public key) for a new position, and the owner is the public key of the issuer LP. An existing position will receive an increase in its balance (or any other modifications on its price range) after verifying that the transaction issuer is indeed the rightful position owner. Mint transactions are initially invoked with a desired amount of token $A$ and token $B$ as input. The underlying AMM algorithms compute the maximum amount of liquidity (based off the input tokens) that the pool can take in at the current moment from both token types. These values represent the share of the pool liquidity now owned by the newly minted or modified position. The committee then deducts the provided liquidity amount (from both token types) from the corresponding LP’s deposit balance.